OpenStudy (anonymous):

Solve P = 2(l + w) for l. What are the missing values in the table? P w l 14 2 ? 22 8 ?

OpenStudy (anonymous):

Just replace the numbers and solve for the missing variable

OpenStudy (anonymous):

P/2=l+w l=P/2-w

OpenStudy (nubeer):

P = 14, w =2 14 = 2( I +2 ) divide both sides by 2 7 = I +2 rest is simple algebra.. same way find the value of I when P = 22 W = 8

OpenStudy (anonymous):

Or a better way is to isolate the variable you are trying to find before plugging in.

OpenStudy (anonymous):

Since we are trying to solve for I we divide both sides by 2 and then subtract W

OpenStudy (anonymous):

P = 2(l + w) \[\frac{p}{2} - w = I\]

OpenStudy (anonymous):

ok thanks